OpenStudy (anonymous):

Gabriel put $6000 in a 2-year CD at 4% interest, compounded monthly. After 2 years, he withdrew all of his money. How much money did he withdrawal?

OpenStudy (anonymous):

Total money = rate * initial investment * time + initial investment T = .04(6000)24 + 6000 T = 240(24) + 6000 T = 5760 + 6000 T = 11760

OpenStudy (anonymous):

the answer either has to be A. $6244.45 B. $6000.00 C. $6254.41 D. $6498.86

OpenStudy (anonymous):

And it's obviously not B.

OpenStudy (anonymous):

The answer is D. I used the wrong formula. Total = initial (1 + rate/n)^(n * time). n is 12, because it's compounded monthly. so T = 6000(1 + .04/12)^(12 * 2) T = 6000(1.0033)^24 T = 6000(1.083143) T= 6498.86

OpenStudy (anonymous):

Thank you for taking the time to show all the steps! A lot more helpful!